OpenStudy (anonymous):

Jessica deposits $5,000 at the end of each year in an account earning 2.45% interest, compounded annually. What is the future value of this annuity after 5 years of investing?

OpenStudy (issy14):

You can solve this by proportions. \[\frac{ 5000 }{ 100 }=\frac{ x }{ 2.5 } = $125 \] She is going to earn $125 dollars on the first year. Since it's compound interest that money will continue to grow and add to the the last year. $125 x 5 = $625 add the total interest to her initial investment 5,000 + 625 =$5625

OpenStudy (issy14):

I'm assuming that she is doing this in different accounts, otherwise, it would be a bit different, first year would be a total of $5125, the second year would be another $5,000. total being $10,125 making a %2.5 to be : $253.12 adding that to the investment being: $10,378 and so on and so forth. It just keeps adding and adding.